# Rapidly-rotating Population III stellar models as a source of primary nitrogen

### v1

@@ 3.1 Transport processes
Figure 1: Internal profiles as a function of the radius and the Lagrangian mass coordinate inside 9 $M_{\odot}$ models half-way through the MS ($X_{c}=0.35$) for different initial rotations: $\upsilon_{\rm ini}/\upsilon_{\rm crit}=$ 0.2 (orange), 0.4 (red) (Murphy et al., 2021a), and 0.7 (green). The grey-shaded zone is the convective core for the moderately-rotating model. Left panel: Radial component of the meridional circulation velocity $U_{r}$. Middle panel: Angular velocity $\Omega$. Right panel: The solid curve show the diffusion coefficient due to shear turbulence $D_{\rm shear}$. The dotted curve shows the effective diffusion coefficient $D_{\rm eff}$ (Y-axis in logarithmic scale and in units of ${\rm cm}^{2}/{\rm s}$). For the purpose of clarity, only the curve for $D_{\rm eff}$ corresponding to $\upsilon_{\rm ini}=0.7\,\upsilon_{\rm crit}$ is shown. In the rest of the radiative zone $D_{\rm eff}$ is in general much smaller than $D_{\rm shear}$.
@@
Figure 1 shows the part of the radial component of the meridional circulation velocity [^4] The radial component of the meridional circulation velocity is given by $u_{r}(r,\theta)=U_{r}(r)P_{2}(\cos\theta)$, where $u_{r}(r,\theta)$ is the radial component of the physical velocity, $\theta$ is the co-latitude, and $P_{2}$ the second Legendre polynomial. $U_{r}$ that depends on the radial coordinate $r$, and the angular velocity in 9 $M_{\odot}$ stellar models half-way through the Main Sequence (MS) phase (e.g. when the mass fraction of hydrogen at the centre $X_{\rm c}=0.35$) for three different initial rotations. Increasing the initial rotation significantly enhances $U_{r}$ both in the radiative envelope (at radial coordinates equal to 0.6-0.7 $R_{\odot}$) and just above the convective core (around 0.2 $R_{\odot}$). In the models sketched in Fig. 1, we can identify at least three changes of sign for $U_{r}$ (when $U_{r}$ is positive the meridional currents transport the angular momentum from the outer regions to the inner ones and the reverse when it is negative). We see that just above the core and further out in the radiative envelope, there are regions where the meridional currents transport the angular momentum inwards, reinforcing there the gradient of the angular velocity $\Omega$. This behaviour cannot be modelled by approximating the angular momentum transport by meridional currents as a diffusive process. Indeed a diffusive process will always tend to smooth any $\Omega-$gradient.
@@
The interactions between the meridional currents, the effects induced by the changes of structure, and the diffusion by shear on the angular momentum distribution result in the variation of the angular velocity with the radius plotted on the middle panel of Fig. 1. It shows that the gradients of $\Omega$ are steeper in the rapidly-rotating models, which implies more chemical mixing by shear. This is consistent with the behaviour of these models in the HRD (see Fig. 5), where indeed the rapidly-rotating models follow, once the chemical mixing had time to be significant, a bluer and more luminous track indicative of a more efficient chemical mixing. The steepness of the $\Omega$-gradient is caused by the core contraction and the transport of the angular momentum. As it has been discussed in many previous papers (see e.g. Meynet & Maeder, 2000), in models without any magnetic fields, angular momentum is mainly transported by meridional currents. In regions where $U_{r}$ is positive, the angular momentum is transported from the outer layers into the inner ones, reinforcing the gradient produced by the core contraction and envelope expansion. As can be seen from the left panel of Fig. 1, $U_{r}$ is positive in a large part of the radiative envelope and get larger values in the faster rotating model making thus the $\Omega$-gradient steeper. Note that such a process cannot be reproduced when the transport of the angular momentum by the meridional currents is described by a diffusive equation as done in some codes.
@@
In the 9 $M_{\odot}$ model, the mixing of the elements is mainly governed by the vertical shear diffusion coefficient ($D_{\rm shear}$), except in a small region above the convective core where the diffusion coefficients describing the results of the action of both the meridional currents and of the horizontal shear turbulence ($D_{\rm eff}$) dominates (see the right panel of Fig. 1). In contrast, in more massive models (typically for the 60 $M_{\odot}$ model and above), the main driver for the transport of the chemical species is $D_{\rm eff}$ (see the lower right panel of Fig. 13). The diffusion coefficient $D_{\rm eff}$ scales with the meridional velocity (Chaboyer & Zahn, 1992), which in turn scales as the inverse of the density. The latter scaling implies that for a given rotation, the meridional velocities are in general larger in the less dense more massive stars. As a numerical example, the maximum value is nearly $8\times 10^{-5}\,{\rm cm}/{\rm s}$ in the rapidly-rotating 9 $M_{\odot}$ model plotted in Fig. 1. In the corresponding 120 $M_{\odot}$ model, the absolute value is of the order of $1\,{\rm cm}/{\rm s}$, hence four orders of magnitude larger (see the upper left panel of Fig. 13). Moreover, it is a negative value indicating that it transports angular momentum from the inner layers to the outer regions. This transport appears more efficient in rapidly-rotating stars.
3.2 Surface He-enrichments
@@
Figure 2: Surface 4He enrichment at the end of the MS phase (see text) as a function of the initial mass for all the Pop III models presented in Table 4.
@@
In Fig. 2 we show the surface 4He enrichment at the end of the MS phase for different initial masses and rotations. The quantity plotted is $\Delta Y=Y_{\rm surf}-Y_{\rm ini}$, where $Y_{\rm ini}=0.2484$ is the initial abundance of 4He. The most rapidly-rotating model (see the green curve) achieves larger surface He enrichment than the less rapidly-rotating ones at the end of the MS phase. Interestingly, the curves for a fixed value of $\upsilon_{\rm ini}/\upsilon_{\rm crit}$ are non-monotonic. They present a concave shape (see the curves for the initial rotation $0.4\,\upsilon_{\rm crit}$ and $0.7\,\upsilon_{\rm crit}$). This shape is due to the following effects: on the one hand, the radius of a lower initial mass star is smaller than the one of a more massive star, implying that, for a given shear diffusion coefficient, the mixing timescale is smaller in the case of a lower initial mass star (the timescale of mixing due to a diffusion coefficient $D$ in a zone of spatial extent $R$ scales as $R^{2}/D$). On the other hand, the total diffusion coefficient that is the sum of $D_{\rm shear}$ and $D_{\rm eff}$ is in general larger in more massive stars compare the right panel of Fig. 1 with the two lower panels of Fig. 13). This makes the mixing timescale shorter over a given distance. These effects, together with the different sizes of the convective cores in the different initial mass models contribute to producing the curves shown in Fig. 2.
3.3 Pop III stars at the critical limit
@@
Figure 3: Evolution of the ratio between the surface angular velocity and the critical one as a function of the fraction of the total MS lifetime ($\tau_{\rm MS}$), for models within the mass range of 9 $M_{\odot}$ $\leqslant M_{\rm ini}\leqslant$ 120 $M_{\odot}$ with different initial rotations. The grey dashed line indicates the critical limit. The moderately-rotating models are from Murphy et al. (2021a).
@@
Figure 3 shows the evolution of the ratio between the surface angular velocity and the critical one during the MS lifetime for models with moderate and fast rotation. The rapidly-rotating models reach the critical velocity at an earlier stage of the evolution during the core H-burning phase (the models with $M_{\rm ini}\geqslant 60\,{M_{\odot}}$ spend more than 50% of their MS lifetime at the critical limit). However, the maximum mass that these model can lose via this process is only at most 2% of their initial mass. So the present models can be considered as evolving at nearly constant mass. We note that it appears plausible that such stars would be surrounded by a decretion disk as Oe-Be stars observed in the present-day Universe. They would be He-rich stars according to the present models.
@@ 3.4 CNO shell boost
Figure 4: Evolutionary track for the 20 $M_{\odot}$ rapidly-rotating model (upper panel) and Kippenhahn diagram (lower panel). The orange and purple points overplotted on the evolutionary track in the upper panel have an internal structure shown in the lower panel at the two times indicated by respectively the vertical orange and purple lines. The blue-shaded zones are the convective areas, while the white zones are the radiative ones. The blue and green curves show the mass coordinates where hydrogen and helium are burning, respectively. Furthermore, the solid curves correspond to the peak of the energy generation rate and the dotted ones to the 10% of the peak energy generation rate for each burning phase.
@@
Ekström et al. (2008) showed that rotational mixing can drive an H-shell boost due to the injection (by rotational mixing and/or convection) of significant amounts of carbon and oxygen from the He-burning core. This boosts the energy generation in the shell and may sometimes give birth to a very extended intermediate convective zone attached to the H-burning shell. The boost derives from the fact that the nuclear reactions of $pp$-chains are taken over by the CNO cycle, with a very different dependence in temperature. In some extreme cases, the H-shell may even become for a time the main source of energy in the star, quenching the energy produced in the He-burning core. When the mixing is gentle and progressive, the star has time to readjust and evolve in a continuous smooth way. When the mixing is much more rapid, it may produce a strong and abrupt change of the structure as the extreme case sketched in Fig. 4. In the following, we shall designate such an event as a CNO shell boost.
@@
The upper panel of Fig. 4 shows the evolutionary track in the HRD of the rapidly-rotating 20 $M_{\odot}$ model. The orange point along the track indicates the end of the core H-burning phase and the beginning of the core He-burning phase. The structure of the star at that stage is given by the vertical orange line in the Kippenhahn diagram shown below. The core He-burning occurs in a convective core with a mass of about 3 $M_{\odot}$, but suddenly the convective core disappears when, due to the process described above, carbon and oxygen are injected into the H-burning shell. This increases the nuclear energy produced in the shell. An intermediate convective shell appears that facilitates still more the feeding of the H-shell in carbon and oxygen. A CNO shell boost occurs. It produces such a strong release of energy that the intermediate convective zone extends from the mass coordinate 2 up to 12 $M_{\odot}$. The purple point along the evolutionary track shows the position in the HRD where the CNO shell boost occurs, and the corresponding internal structure is indicated by the vertical purple line in the Kippenhahn diagram. From that stage onwards, the He-burning convective core disappears for a while. The star is mainly powered by H-burning (in a very extended shell). This produces a “new” MS track in the HRD.
@@
A similar figure for the 60 $M_{\odot}$ is shown in Fig. 14. Qualitatively, we have a similar evolution as the 20 $M_{\odot}$ model. After a phase of core He-burning (between time coordinates around 5.52 and 5.32 in logarithm of the remaining time in years) during which no convective shell is associated with the H-burning shell, an intermediate convective shell appears above the H-burning region. This never succeeds in quenching the nuclear reaction in the core but significantly reduces the size of the He-burning convective core. Such a feature in very massive stars may prevent the star from entering the pair-instability regime and may have importance for determining the lower limit of the black hole mass gap (Farrell et al., 2021).
4 The rapidly-rotating grid
@@
Figure 5: HRD of the Pop III models from Murphy et al. (2021a) (blue: $\upsilon_{\rm ini}=0$, red: $\upsilon_{\rm ini}=0.4\,\upsilon_{\rm crit}$) and the rapidly-rotating models computed for this work (green: $\upsilon_{\rm ini}=0.7\,\upsilon_{\rm crit}$). The symbols drawn at the end of the tracks mark the final evolutionary stage attained: a star for models that have completed the core He-burning phase, a circle for models that are either during the core C-burning phase or at the end of it, and a diamond for models that are after the core C-burning phase. The models without any symbol were stopped in the core He-burning phase. Iso-radius lines are drawn in dotted grey. The pink arrows indicate the transition between the pp-contraction phase and the pp-CNO-sustained phase.
@@
Table 1: Properties of Pop III models at the end of the core H- and He-burning phases with different initial surface velocities. Models with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=$ 0, 0.2, and 0.4 are from Murphy et al. (2021a). Models that have not finished the core He-burning phase are indicated with an horizontal dash.
@@
As is well known (see e.g. Cassisi & Castellani, 1993; Marigo et al., 2001; Ekström et al., 2008) the absence of any CNO elements in Pop III stars leaves only $pp$-chains at the beginning for producing nuclear energy. The energy released by this process is however not enough to compensate the losses at the surface and the star contracts (we call here this phase the pp-contraction phase). This occurs until temperatures are high enough to activate some synthesis of carbon by triple $\alpha$ reactions. From this stage on, the star reaches a radiative equilibrium in which the energy losses at the surface can be compensated by the energy produced at the centre through nuclear reactions involving pp and CNO reactions (pp-CNO-sustained phase).
@@
In the evolutionary tracks presented in Fig. 5, the transition between the pp-contraction phase and the pp-CNO-sustained phase is clearly visible in the case of the 9 and 20 $M_{\odot}$ models. It corresponds to the points when the effective temperature begins to decrease as the luminosity increases (pink arrows on Fig. 5). In more massive stars, the pp-contraction phase is too short to be visible.
@@
In the upper mass range, the models show inflexion points (well visible for the $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.7$ curves, less so for the slower-rotating models) during the MS phase. These inflexion points occur when the surface velocity reaches the critical limit. The mechanical mass loss slows down the surface below the critical value but its secular evolution will bring it back to the critical limit. This back-and-forth evolution produces the oscillations seen in effective temperatures along those tracks [^5] During the short timescale of mechanical mass loss, the luminosity remains constant. Since when losing mass, the star loses angular momentum and becomes less oblate, for example, less deformed by rotation, its surface area, $S$, decreases and hence, through the relation $L=S\sigma T_{\rm eff}^{4}$, its effective temperature increases..
@@
Our rapidly-rotating models are slightly shifted to the red in the first part of the MS phase. Then in the second part, they cross the slower-rotating track, reaching a higher luminosity. Except in the case of the 9 $M_{\odot}$, the rapidly-rotating tracks show a turn-off at the end of the MS phase at lower effective temperatures than their slower-rotating counterparts. These two trends are well-known effects of rotation: the first one shows the hydrostatic effects of rotation (see e.g. Faulkner et al., 1968; Kippenhahn et al., 1970; Meynet & Maeder, 1997). The second one shows the mixing effects of rotation, mainly the enrichment in He of the radiative envelope (Langer, 1992; Meynet & Maeder, 2000, see also the discussion in Sect. 3.2) and the impact of the mixing on the convective core mass evolution (Talon et al., 1997; Deupree, 1998; Heger & Langer, 2000).
@@
From Fig. 5, we can see how changing the initial rotation impacts the nature of the star at a given evolutionary stage. As an example considering the 20 $M_{\odot}$ model, the non-rotating model after the end of the core C-burning stage is a Blue Supergiant (BSG), while the rapidly-rotating one is a Red Supergiant (RSG). We may thus expect very different supernova (SN) light curves from the non-rotating and rapidly-rotating models in case an explosion occurs at the time of the final core collapse (see e.g. Yoon et al., 2012; Aryan et al., 2023) (note that we do not expect any significant change of the surface conditions in the few tens of years that remain after the C-burning phase for the star to reach the final core-collapse phase).
@@
Although the initial angular momentum of those models is high, the mixing is not efficient enough for making them evolve homogeneously (Maeder, 1987). Other rotating models (see e.g. Szécsi et al., 2015) show a chemical homogeneous evolution for even slower initial rotation for masses above 26 $M_{\odot}$ at $Z=0.0002$. Here of course we have metal-free Pop III stars, which may explain in part this difference. Besides, the physics of rotation is different between the present models and those e.g. of Szécsi et al. (2015). The latter include the Tayler-Spruit dynamo (Spruit, 2002; Heger et al., 2005). We would like to remind the reader of the fact that the term “rotating models” may actually correspond to very different physics and also to very different outputs. The results presented here correspond to one choice (non-ganetic models with specific expressions for the diffusion coefficients). The impact of other choices is explored in a dedicated work (Nandal et al., 2023b). It is still difficult to determine which kind of physics for the transport processes due to rotation are the most favoured. At solar metallicity, models with a more efficient angular momentum transport than those considered in this work are favoured. Indeed, transport accounting for the Tayler-Spruit dynamo produces much better agreement with asteroseismic data of subgiants (Moyano et al., 2023) [^6] Note that these models include what has been called the Tayler-Spruit calibrated dynamo theory proposed by Eggenberger et al. (2022). This implementation is different from the one accounted for in the models by Szécsi et al. (2015). than non-magnetic models as those presented here. However, we are exploring here a very special regime of initial composition (primordial stars) where it is not guaranteed that magnetic fields play a similar role. Also the transport of the chemical species remains to be studied in more details. Its efficiency (and thus its capability of driving a homogeneous evolution) depends mainly on the expression used for the horizontal shear turbulent coefficient. Thus, unfortunately without more dedicated study, it is still difficult to decide which type of models are the most representative of the bulk Pop III star populations. We just retain here that the conditions for obtaining a homogeneous evolution heavily depends on the physics assumed for both the transport of the chemical elements and of the angular momentum.
@@
Some features of the Pop III stellar models at the end of the core H- and He-burning phase are presented in Table 4. Columns $1-2$ give the initial mass and the initial rotation on the ZAMS. Columns $3-6$ show respectively the duration of the MS lifetime, and the values at the end of the core H-burning phase of the ratios between the surface to the critical velocity, the linear surface velocities at the equator, and the mass fractions of helium at the surface. The same quantities for the core He-burning phase are indicated in columns $7-10$.

### v2

@@ 3.1 Transport processes
Figure 1: Internal profiles as a function of the radius and the Lagrangian mass coordinate inside 9 $M_{\odot}$ models halfway through the MS ($X_{c}=0.35$) for different initial rotations: $\upsilon_{\rm ini}/\upsilon_{\rm crit}=$ 0.2 (orange), 0.4 (red) (Murphy et al., 2021a), and 0.7 (green). The grey-shaded zone is the convective core of the moderately rotating model. Left panel: Radial component of the meridional circulation velocity, $U_{r}$. Middle panel: Angular velocity, $\Omega$. Right panel: The solid curve shows the diffusion coefficient due to shear turbulence, $D_{\rm shear}$. The dotted curve shows the effective diffusion coefficient, $D_{\rm eff}$ (y-axis in logarithmic scale and in units of ${\rm cm}^{2}/{\rm s}$). For clarity, only the curve for $D_{\rm eff}$ corresponding to $\upsilon_{\rm ini}=0.7\,\upsilon_{\rm crit}$ is shown. In the rest of the radiative zone, $D_{\rm eff}$ is in general much smaller than $D_{\rm shear}$.
@@
Figure 1 shows the part of the radial component of the meridional circulation velocity [^4] The radial component of the meridional circulation velocity is given by $u_{r}(r,\theta)=U_{r}(r)P_{2}(\cos\theta)$, where $u_{r}(r,\theta)$ is the radial component of the physical velocity, $\theta$ is the co-latitude, and $P_{2}$ the second Legendre polynomial., $U_{r}$, that depends on the radial coordinate, $r$, and the angular velocity in 9 $M_{\odot}$ stellar models halfway through the MS phase (e.g. when the mass fraction of hydrogen at the centre is $X_{\rm c}=0.35$) for three different initial rotations. Increasing the initial rotation significantly enhances $U_{r}$ both in the radiative envelope (at radial coordinates equal to 0.6-0.7 $R_{\odot}$) and just above the convective core (at around 0.2 $R_{\odot}$). In the models sketched in Fig. 1, we can identify at least three changes of sign for $U_{r}$ (when $U_{r}$ is positive, the meridional currents transport the angular momentum from the outer regions to the inner ones, and the reverse is true when it is negative). We see that just above the core, and further out in the radiative envelope, there are regions where the meridional currents transport the angular momentum inwards, reinforcing there the gradient of the angular velocity, $\Omega$. This behaviour cannot be modelled by approximating the angular momentum transport by meridional currents as a diffusive process. Indeed, a diffusive process will always tend to smooth any $\Omega-$gradient.
@@
The interactions between the meridional currents, the effects induced by the changes of structure, and the diffusion by shear on the angular momentum distribution result in the variation of the angular velocity with the radius plotted in the middle panel of Fig. 1. It shows that the gradients of $\Omega$ are steeper in the rapidly rotating models, which implies more chemical mixing by shear. This is consistent with the behaviour of these models in the HRD (see Fig. 5), where indeed the rapidly rotating models follow, once the chemical mixing has had time to be significant, a bluer and more luminous track that is indicative of a more efficient chemical mixing. The steepness of the $\Omega-$gradient is caused by the core contraction and the transport of the angular momentum. As has been discussed in many previous papers (see e.g. Meynet & Maeder, 2000), in models without any magnetic fields, angular momentum is mainly transported by meridional currents. In regions where $U_{r}$ is positive, the angular momentum is transported from the outer layers into the inner ones, reinforcing the gradient produced by the core contraction and envelope expansion. As can be seen from the left panel of Fig. 1, $U_{r}$ is positive in a large part of the radiative envelope and reaches larger values in the faster rotating model, thus making the $\Omega-$gradient steeper. Such a process cannot be reproduced when the transport of the angular momentum by the meridional currents is described by a diffusive equation, as is done in some codes.
@@
In the 9 $M_{\odot}$ model, the mixing of the elements is mainly governed by the vertical shear diffusion coefficient ($D_{\rm shear}$), except in a small region above the convective core where the diffusion coefficients describing the results of the action of both the meridional currents and of the horizontal shear turbulence ($D_{\rm eff}$) dominate (see the right panel of Fig. 1). In contrast, in more massive models (typically for the 60 $M_{\odot}$ model and above), the main driver for the transport of the chemical species is $D_{\rm eff}$ (see the lower right panel of Fig. 13). The diffusion coefficient, $D_{\rm eff}$, scales with the meridional velocity (Chaboyer & Zahn, 1992), which in turn scales as the inverse of the density. The latter scaling implies that for a given rotation the meridional velocities are in general larger in less dense, more massive stars. As a numerical example, the maximum value is nearly $8\times 10^{-5}\,{\rm cm}/{\rm s}$ in the rapidly rotating 9 $M_{\odot}$ model plotted in Fig. 1. In the corresponding 120 $M_{\odot}$ model, the absolute value is of the order of $1\,{\rm cm}/{\rm s}$, and hence four orders of magnitude larger (see the upper left panel of Fig. 13). Moreover, it is a negative value, indicating that it transports angular momentum from the inner layers to the outer regions. This transport appears to be more efficient in rapidly rotating stars.
3.2 Surface He-enrichment
@@
Figure 2: Surface 4He-enrichment at the end of the MS phase (see text) as a function of the initial mass for all the Pop III models presented in Table 4.
@@
In Fig. 2 we show the surface 4He-enrichment at the end of the MS phase for different initial masses and rotations. The quantity plotted is $\Delta Y=Y_{\rm surf}-Y_{\rm ini}$, where $Y_{\rm ini}=0.2484$ is the initial abundance of 4He. The most rapidly rotating model (see the green curve) achieves larger surface He-enrichment than the less rapidly rotating ones at the end of the MS phase. Interestingly, the curves for a fixed value of $\upsilon_{\rm ini}/\upsilon_{\rm crit}$ are non-monotonic. They present a concave shape (see the curves for the initial rotation, $0.4\,\upsilon_{\rm crit}$ and $0.7\,\upsilon_{\rm crit}$). This shape is due to the following effects: on the one hand, the radius of a star with a lower initial mass is smaller than the one of a more massive star, implying that, for a given shear diffusion coefficient, the mixing timescale is smaller in the case of a lower initial mass star (the timescale of mixing due to a diffusion coefficient, $D$, in a zone of spatial extent, $R$, scales as $R^{2}/D$). On the other hand, the total diffusion coefficient that is the sum of $D_{\rm shear}$ and $D_{\rm eff}$ is in general larger in more massive stars (compare the right panel of Fig. 1 with the two lower panels of Fig. 13). This makes the mixing timescale shorter over a given distance. These effects, together with the different sizes of the convective cores in the different initial mass models, contribute to producing the curves shown in Fig. 2.
3.3 Population III stars at the critical limit
@@
Figure 3: Evolution of the ratio between the surface angular velocity and the critical one as a function of the fraction of the total MS lifetime ($\tau_{\rm MS}$), for models within the mass range of 9 $M_{\odot}$ $\leqslant M_{\rm ini}\leqslant$ 120 $M_{\odot}$ with different initial rotations. The dashed grey line indicates the critical limit. The moderately rotating models are from Murphy et al. (2021a).
@@
Figure 3 shows the evolution of the ratio between the surface angular velocity and the critical one during the MS lifetime for models with moderate and fast rotation. The rapidly rotating models reach the critical velocity at an earlier stage of the evolution, during the core H-burning phase (the models with $M_{\rm ini}\geqslant 60\,{M_{\odot}}$ spend more than 50% of their MS lifetime at the critical limit). However, the maximum mass that these models can lose via this process is only at most 2% of their initial mass, so the present models can be considered to be evolving at a nearly constant mass. We note that it appears plausible that such stars would be surrounded by a decretion disc as Oe-Be stars observed in the present-day Universe. They would be He-rich stars according to the present models.
@@ 3.4 CNO shell boost
Figure 4: Evolutionary track for the 20 $M_{\odot}$ rapidly rotating model (upper panel) and Kippenhahn diagram (lower panel). The orange and purple points overplotted on the evolutionary track in the upper panel have an internal structure shown in the lower panel at the two times indicated by the vertical orange and purple lines, respectively. The blue-shaded zones are the convective areas, while the white zones are the radiative ones. The blue and green curves show the mass coordinates where hydrogen and helium are burning, respectively. Furthermore, the solid curves correspond to the peak of the energy generation rate and the dashed ones to 10% of the peak energy generation rate for each burning phase.
@@
Ekström et al. (2008) showed that rotational mixing can drive an H-shell boost due to the injection (by rotational mixing and/or convection) of significant amounts of carbon and oxygen from the He-burning core. This boosts the energy generation in the shell and may sometimes give birth to a very extended intermediate convective zone attached to the H-burning shell. The boost derives from the fact that the nuclear reactions of $pp$-chains are taken over by the CNO cycle, with a very different dependence on temperature. In some extreme cases, the H-shell may even become for a time the main source of energy in the star, quenching the energy produced in the He-burning core. When the mixing is gentle and progressive, the star has time to readjust and evolve in a continuous, smooth way. When the mixing is much more rapid, it may produce a strong and abrupt change of the structure like the extreme case sketched in Fig. 4. In the following, we designate such an event a CNO shell boost.
@@
The upper panel of Fig. 4 shows the evolutionary track in the HRD of the rapidly rotating 20 $M_{\odot}$ model. The orange point along the track indicates the end of the core H-burning phase and the beginning of the core He-burning phase. The structure of the star at that stage is given by the vertical orange line in the Kippenhahn diagram shown below. The core He-burning occurs in a convective core with a mass of about 3 $M_{\odot}$, but suddenly the convective core disappears when, due to the process described above, carbon and oxygen are injected into the H-burning shell. This increases the nuclear energy produced in the shell. An intermediate convective shell appears that facilitates still more the feeding of the H-shell in carbon and oxygen. A CNO shell boost occurs. It produces such a strong release of energy that the intermediate convective zone extends from the mass coordinate 2 up to 12 $M_{\odot}$. The purple point along the evolutionary track shows the position in the HRD where the CNO shell boost occurs, and the corresponding internal structure is indicated by the vertical purple line in the Kippenhahn diagram. From that stage onwards, the He-burning convective core disappears for a while. The star is mainly powered by H-burning (in a very extended shell). This produces a ‘new’ MS track in the HRD.
@@
A similar figure for the 60 $M_{\odot}$ is shown in Fig. 14. Qualitatively, we have a similar evolution as the 20 $M_{\odot}$ model. After a phase of core He-burning (between time coordinates around 5.52 and 5.32 in a logarithm of the remaining time in years) during which no convective shell is associated with the H-burning shell, an intermediate convective shell appears above the H-burning region. This never succeeds in quenching the nuclear reaction in the core but significantly reduces the size of the He-burning convective core. Such a feature in very massive stars may prevent the star from entering the pair-instability regime and may have importance in determining the lower limit of the black hole mass gap (Farrell et al., 2021).
4 The rapidly rotating grid
@@
Figure 5: HRD of the Pop III models from Murphy et al. (2021a) (blue: $\upsilon_{\rm ini}=0$, red: $\upsilon_{\rm ini}=0.4\,\upsilon_{\rm crit}$) and the rapidly rotating models computed for this work (green: $\upsilon_{\rm ini}=0.7\,\upsilon_{\rm crit}$). The symbols drawn at the end of the tracks mark the final evolutionary stage attained: a star for models that have completed the core He-burning phase, a circle for models that are either in the core C-burning phase or at the end of it, and a diamond for models that are after the core C-burning phase. The models without any symbol were stopped in the core He-burning phase. Iso-radius lines are drawn in dotted grey. The pink arrows indicate the transition between the pp-contraction phase and the pp-CNO-sustained phase.
@@
Table 1: Properties of Pop III models at the end of the core H- and He-burning phases with different initial surface velocities. Models with $\upsilon_{\rm ini}/\upsilon_{\rm crit}=$ 0, 0.2, and 0.4 are from Murphy et al. (2021a). Models that have not finished the core He-burning phase are indicated with a horizontal dash.
@@
As is well known (see e.g. Cassisi & Castellani, 1993; Marigo et al., 2001; Ekström et al., 2008), the absence of any CNO elements in Pop III stars leaves only $pp$-chains at the beginning to produce nuclear energy. The energy released by this process is however not enough to compensate for the losses at the surface and the star contracts (here we call this phase the pp-contraction phase). This occurs until temperatures are high enough to activate some synthesis of carbon by triple $\alpha$ reactions. From this stage on, the star reaches a radiative equilibrium in which the energy losses at the surface can be compensated for by the energy produced at the centre through nuclear reactions involving pp and CNO reactions (pp-CNO-sustained phase).
@@
In the evolutionary tracks presented in Fig. 5, the transition between the pp-contraction phase and the pp-CNO-sustained phase is clearly visible in the case of the 9 and 20 $M_{\odot}$ models. It corresponds to the points when the effective temperature begins to decrease as the luminosity increases (pink arrows in Fig. 5). In more massive stars, the pp-contraction phase is too short to be visible.
@@
In the upper mass range, the models show inflexion points (clearly visible for the $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.7$ curves, less so for the slower-rotating models) during the MS phase. These inflexion points occur when the surface velocity reaches the critical limit. The mechanical mass loss slows down the surface below the critical value but its secular evolution will bring it back to the critical limit. This back-and-forth evolution produces the oscillations seen in effective temperatures along those tracks [^5] During the short timescale of mechanical mass loss, the luminosity remains constant. Since when losing mass, the star loses angular momentum and becomes less oblate, for example, less deformed by rotation, its surface area, $S$, decreases and hence, through the relation $L=S\sigma T_{\rm eff}^{4}$, its effective temperature increases..
@@
Our rapidly rotating models are slightly shifted to the red in the first part of the MS phase. Then, in the second part, they cross the slower-rotating track, reaching a higher luminosity. Except in the case of the 9 $M_{\odot}$, the rapidly rotating tracks show a turn-off at the end of the MS phase at lower effective temperatures than their slower-rotating counterparts. These two trends are well-known effects of rotation: the first one shows the hydrostatic effects of rotation (see e.g. Faulkner et al., 1968; Kippenhahn et al., 1970; Meynet & Maeder, 1997); the second one shows the mixing effects of rotation, mainly the enrichment in He of the radiative envelope (Langer, 1992; Meynet & Maeder, 2000, see also the discussion in Sect. 3.2) and the impact of the mixing on the convective core mass evolution (Talon et al., 1997; Deupree, 1998; Heger & Langer, 2000).
@@
From Fig. 5, we can see how changing the initial rotation impacts the nature of the star at a given evolutionary stage. As an example considering the 20 $M_{\odot}$ model, the non-rotating model after the end of the core C-burning stage is a blue supergiant (BSG), while the rapidly rotating one is a red supergiant (RSG). We may thus expect very different SN light curves from the non-rotating and rapidly rotating models in case an explosion occurs at the time of the final core collapse (see e.g. Yoon et al., 2012; Aryan et al., 2023) (it should be noted that we do not expect any significant change in the surface conditions in the few tens of years that remain after the C-burning phase for the star to reach the final core-collapse phase).
@@
Although the initial angular momentum of those models is high, the mixing is not efficient enough to make them evolve homogeneously (Maeder, 1987). Other rotating models (see e.g. Szécsi et al., 2015) show a chemically homogeneous evolution for an even slower initial rotation for masses above 26 $M_{\odot}$ at $Z=0.0002$. Here of course we have metal-free Pop III stars, which may explain in part this difference. Besides, the physics of rotation is different between the present models and those of, for example, Szécsi et al. (2015). The latter include the Tayler-Spruit dynamo (Spruit, 2002; Heger et al., 2005). We would like to remind the reader of the fact that the term ‘rotating models’ may actually correspond to very different physics and also to very different outputs. The results presented here correspond to one choice (non-magnetic models with specific expressions for the diffusion coefficients). The impact of other choices is explored in a dedicated work (Nandal et al., 2024). It is still difficult to determine which kind of physics for the transport processes due to rotation are the most favoured. At solar metallicity, models with a more efficient angular momentum transport than those considered in this work are favoured. Indeed, transport accounting for the Tayler-Spruit dynamo produces a much better agreement with asteroseismic data of subgiants (Moyano et al., 2023) [^6] Note that these models include what has been called the Tayler-Spruit calibrated dynamo theory proposed by Eggenberger et al. (2022). This implementation is different from the one accounted for in the models by Szécsi et al. (2015). than non-magnetic models like those presented here. However, we are exploring here a very special regime of initial composition (primordial stars) where it is not guaranteed that magnetic fields play a similar role. Also, the transport of the chemical species remains to be studied in more detail. Its efficiency (and thus its capability of driving a homogeneous evolution) depends mainly on the expression used for the horizontal shear turbulent coefficient. Thus, without more dedicated study, it is unfortunately still difficult to decide which types of models are the most representative of the bulk Pop III star populations. We just retain here that the conditions for obtaining a homogeneous evolution heavily depend on the physics assumed for the transport of both the chemical elements and the angular momentum.
@@
Some features of the Pop III stellar models at the end of the core H- and He-burning phase are presented in Table 4. Columns $1-2$ give the initial mass and the initial rotation on the ZAMS. Columns $3-6$ show, respectively, the duration of the MS lifetime, the values at the end of the core H-burning phase of the ratios between the surface to the critical velocity, the linear surface velocities at the equator, and the mass fractions of helium at the surface. The same quantities for the core He-burning phase are indicated in columns $7-10$.